# Unified Smooth Vector Graphics: Modeling Gradient Meshes and Curve-based Approaches Jointly as Poisson Problem

### v1

@@ 6.3 Implementation Detail
The turning number in Eq. (11) is calculated from the polyline discretization, previously described in Section 5.3, by summing up the discrete turning angles. A loop is nested inside another loop if all its discrete polyline vertices are contained. To determine containment bounding boxes are used to find candidates and a detailed check is performed by using the winding angle theorem (Jacobson et al., 2013; Spainhour et al., 2024), which is a numerically robust way of testing containment.
@@ 7.1 Definition
with the patch Laplacian $\mathbf{f}(\mathbf{x})$ as given by Eq. (12) and subject to the patch boundary conditions $\mathcal{B}$} that arise from the patch boundary curves on $\partial\Omega$, as introduced in Eqs. (7)–(8). There are many ways how the Poisson problem could be solved, including direct solvers, iterative solvers, or Monte Carlo methods. We refer to a survey (Tian and Günther, 2023) for an overview of solvers in the smooth vector graphics literature and also point out several recent solvers, which appeared after the survey (Bang et al., 2023; Sawhney et al., 2023; Sugimoto et al., 2023; Yu et al., 2024; Chen et al., 2024). Since the PDE solver is orthogonal to our contributions, we utilized a simple Jacobi relaxation for simplicity (Jeschke et al., 2009):
@@ 7.2 Implementation Detail
    interior pixel ($k_{i,j}=\textrm{interior}$), then the weights of all adjacent pixels are set to 11, i.e., $w_{i+1,j}=w_{i-1,j}=w_{i,j+1}=w_{i,j-1}=1$.
@@
    boundary pixel with Dirichlet condition ($k_{i,j}=\textrm{Dirichlet}$), then Eq. (14) is not executed. Instead, the color of the pixel is already determined by the boundary condition.
@@
    boundary pixel with Neumann condition ($k_{i,j}=\textrm{Neumann}$), then the weight of the adjacent pixel $w_{i+1,j}$ is set to 11 (otherwise 0) if it is non-separated by a patch boundary curve $s^{h}_{i,j}=\textrm{open}$ (analogous for the other neighbor pixels). This prevents color leaking over curves with Neumann condition that are contained inside a patch, see for example Fig. 5.
@@ 8 Results
In the following section, we present our results using scenes that contain gradient meshes, diffusion curves, and Poisson curves. We made all test scenes available online (Tian and Günther, 2024). All images are rendered at a resolution of $1024\times 1024$.
(a) input meshes
(b) input curves
@@ 8.1 Qualitative Results
We begin with a description of the test scenes. For all test scenes, the input meshes and curves, the undirected edge graph, our unified patch representation, and the final rendering result are shown. To illustrate the input primitives, we depict the gradient mesh control polygon with solid lines, where different colors are applied for each mesh. Input boundary curves are shown as double-sided curve with an illustration of the respective boundary condition on the left and right. Dirichlet conditions directly show the corresponding color, while Neumann conditions are illustrated with a dash-dot pattern. Poisson curves are visualized with a long dash pattern. The undirected edge graph is visualized by solid black curves and the vertices are depicted with black circles. Note that Poisson curves are not part of the edge graph, since they are rendered additively onto the target Laplacian function. The unified patch representation gives all pixels that belong to the same region a common color. The colors are assigned from a custom color table such that the regions are visually distinguishable. The color is not related to the image content. The last column shows the final result.
@@
Inspired from an ancient Chinese painting (the Auspicious Cranes by Emperor Huizong), we modeled a crane with diffusion curves, which is shown in Fig. 1. The boundary of the crane is largely modeled with Neumann conditions. We modeled the clouds with two $3\times 4$ gradient meshes. In addition, we added diffusion curves and Poisson curves on the gradient meshes for the layered and detailed effects of clouds. The background is formed from a $3\times 3$ gradient mesh that mimics the similar color transition of the inspirational image. The image was rendered using 1 k Jacobi iterations.
@@
Fig. 12 (first row) shows an extension of the poivron scene from Orzan et al. (2008). We replaced the background with a $30\times 30$ gradient mesh that was fitted with Adobe Illustrator to a photograph inside a restaurant. The gradient mesh’s smoothness creates the impression of defocus blur, putting emphasis on the pepper in the foreground. We took a bite out of the pepper by adding further diffusion curves. The outer boundary of the pepper is modeled with homogeneous Neumann conditions, such that it can be placed in the scene without causing color diffusion into the background. The shadow on the table is modeled with Poisson curves. The background contains many color gradients, which would have been cumbersome to model with diffusion curves only. To render the image we performed 10 k Jacobi iterations using Eq. (14).
@@
Fig. 12 (second row) shows a colorful sunset. The color gradient in the background contains multiple hues and is modeled with a $10\times 3$ gradient mesh. We also deformed the spatial control points for artistic effects. Since it is interrupted by other curve primitives in-between, the color gradient would have been difficult to model with diffusion curves only. The sun is drawn using a diffusion curve, which allows for the modeling of the faint glow due to scattering. The sun diffuses its light into the background. On the top and bottom of the image, we placed diffusion curves that contrast the nature background with expressive wavy shapes. This image was rendered using 15 k Jacobi iterations.
@@
In Fig. 12 (third row), we modeled two soap bubbles. We used a $1\times 3$ gradient mesh to model the background sky for smooth color transitions. The bubbles are created by two $4\times 4$ gradient meshes. The smooth color transitions are difficult to model directly with diffusion curves. This example shows that a vivid scene can be represented by a small number of primitives when combining mesh-based and curve-based methods. We rendered the image with 15 k Jacobi iterations.
@@
Table 1: Performance measurements for the various test scenes used in the paper. The columns list the corresponding figure, the number of input diffusion curves (#DC), the number of input Poisson curves (#PC), the number of input gradient meshes (#GM) for which we list the grid resolution in brackets, the number of vertices (#V) and edges (#E) in the edge graph, the number of patches (#P), the computation time to form the edge graph, the computation time to traverse the edge graph, and the time needed to resolve the nesting of patches. All timings are in milliseconds.
Scene Figure #DC #PC #GM #V #E #P Construct graph Trav. graph Create patches
Crane Fig. 1 68 5 3 160 124 8 16 5 13
Pepper Fig. 12 (first row) 36 3 1 184 165 8 11 9 12
Sunset Fig. 12 (second row) 38 6 1 88 108 37 16 2 7
Bubble Fig. 12 (third row) 9 0 3 62 53 3 2 1 1
Ladybug Fig. 12 (fourth row) 59 8 2 111 113 20 14 2 16
Portal Fig. 12 (fifth row) 87 1 3 211 167 17 20 3 14
Ladybug
The test scene in Fig. 12 (fourth row) is based on the ladybug diffusion curve model of Orzan et al. (2008). We made minor adjustments to the original curve endpoints to remove edge crossings that would otherwise cause color leakage. In addition, we set the outer boundary of the ladybug to Neumann condition. We placed the ladybug in a beach scene, which contains an ocean and a blue sky. The background is modeled using a $11\times 3$ gradient mesh. In addition, we model a leaf on which the ladybug sits with a $6\times 2$ gradient mesh, which captures the subtle color changes without smoothing. The leaf veins are created with Poisson curves, and the outer boundary is modeled with diffusion curves to add a scattering effect. We rendered the image using 2 k Jacobi iterations.
Portal
The last test scene in Fig. 12 (fifth row) shows a portal. We created the portal with three gradient meshes: a $4\times 3$ background mesh, and two $3\times 4$ gradient meshes for the outer and internal parts of the portal. The rendered portal contains smooth color transitions to represent light and shadows. Inside the portal, We modified the zephyr diffusion curve model of Orzan et al. (2008). The diffusion curves diffuse color into the portal. The mesh Laplacian provides richer effects and more potential for artistic control compared to curve-only results. The image was rendered using 15 k Jacobi iterations.
@@ 8.2 User Interaction
When individual input primitives are moved in the scene by the user during editing, intersections with other primitives will occur that lead to the creation of new patches or the disappearance of other patches. To ensure a consistent editing experience, we made sure that all user input is specified on the extended gradient meshes, extended diffusion curves, and Poisson curves alone. On the front end, the user is not accessing the edge graph or the patch representation, which are both computed fully automatically from the input primitives. Fig. 13 shows the rendered output when taking a closed diffusion curve and moving it across a scene. Intersections with other diffusion curves and gradient meshes are created, which causes new patches to form or disappear. Using our data structures, the rendering results remain consistent throughout the editing process.
@@ 8.3 Performance Analysis
We evaluated our method on an Intel Core i9-10980XE with 4.6 Ghz and an Nvidia RTX 2080 TI GPU. The performance measurements are reported in Table 1. To give context we report for each scene statistics about the complexity of the input data (number of diffusion curves, number of Poisson curves, number of gradient meshes), statistics on the complexity of the edge graph (number of vertices and edges), and the number of patches in our patch representation. We report the time it takes to compute the edge graph, the time it takes to traverse the edge graph during the patch construction, and the time it takes to resolve the nesting relationships.
Across all test scenes, the edge graph construction was at around 12-17 ms, the edge graph traversal at 4-10 ms, and the patch creation at 13-45 ms. In sum, we achieved across all scenes an interactive frame rate (29-67 milliseconds per frame). Resolving the nested patches is currently the bottleneck, since it requires the traversals to determine the clockwise/anti-clockwise turn numbers in Eq. (11) and the test whether patches are fully contained within each other. While we incorporated bounding volume hierarchies to accelerate the tests, we think that further improvements for the future could include partial updates of both the edge graph and the patch representation whenever an object moves. Further, a hierarchical discretization could be used to accelerate the winding angle tests. Despite us rebuilding the graph and patches from scratch, we still obtained interactive results.
@@ 8.4 Discussion
The edge graph and our resulting patch construction directly depend on the quality of the input primitives. If the input primitives contain tiny gaps or small crossings, then a corresponding patch representation would carry on those artifacts. In recent years, more exact vector graphics rendering algorithms have been developed and utilized, such as Walk-on-Spheres (Sawhney and Crane, 2020), Walk-on-Boundaries (Sugimoto et al., 2023), or Walk-on-Stars (Sawhney et al., 2023). These approaches likewise suffer from inaccuracies in the scene description, since they are able to resolve the color transport through narrow gaps. More research towards automatic input cleaning or user interfaces that naturally steer the user towards more accurate scene descriptions would be helpful for both scene modeling and rendering.
@@ 9 Conclusions
In this paper, we unified the mathematical modeling of mesh-based vector graphics and curve-based vector graphics, which allows for the first time to include both types of primitives in the same scene. For this, we rephrased the interpolation inside gradient meshes as solution to a Poisson problem by searching for a color field that matches the Laplacian of the gradient mesh. We developed a four-stage pipeline that takes gradient meshes, diffusion curves, and Poisson curves as input. First, we extended gradient meshes and diffusion curves by enabling the specification of Neumann conditions in addition to the usual Dirichlet conditions. Second, our approach handles arbitrary intersections of diffusion curves and gradient meshes, since all intersections are resolved through the construction of an edge graph. Third, from the edge graph, non-overlapping patches with well-defined boundary conditions and a target Laplacian are derived. Fourth, a standard Poisson problem is solved in the interior of each patch to synthesize the output image. Our pipeline is free of parameters and the final rasterized images can be computed with any off-the-shelf Poisson solver.
In the past, research on mesh-based and curve-based smooth vector graphics followed independent research threads, which concentrated on the editing, rasterization, or vectorization of the individual primitives. We hope that the unified treatment will spur further research that promotes synergies of the two approaches, for example regarding novel image synthesis or vectorization methods. The next step in the generalization will include the incorporation of bi-harmonic diffusion curves (Finch et al., 2011), for which harmonic solutions can be searched in the Laplacian domain. Similar to common content creation tools, it could be useful to introduce multiple layers of smooth vector graphics.

### v2

@@ 6.3 Implementation Detail
The turning number in Eq. (11) is calculated from the polyline discretization, previously described in Section 5.3, by summing up the discrete turning angles. A loop is nested inside another loop if one of its vertices is contained in the other loop, since loops are by construction intersection free. To determine containment bounding boxes are used to find candidates and a detailed check is performed by using the winding angle theorem (Jacobson et al., 2013; Spainhour et al., 2024), which is a numerically robust way of testing containment.
@@ 7.1 Definition
with the patch Laplacian $\mathbf{f}(\mathbf{x})$ as given by Eq. (12) and subject to the patch boundary conditions $\mathcal{B}$} that arise from the patch boundary curves on $\partial\Omega$, as introduced in Eqs. (7)–(8). There are many ways how the Poisson problem could be solved, including direct solvers, iterative solvers, or Monte Carlo methods. We refer to a survey (Tian and Günther, 2024) for an overview of solvers in the smooth vector graphics literature and also point out several recent solvers, which appeared after the survey (Bang et al., 2023; Sawhney et al., 2023; Sugimoto et al., 2023; Miller et al., 2024b; Chen et al., 2024). Since the PDE solver is orthogonal to our contributions, we utilized a simple (multi-grid) Jacobi relaxation (Jeschke et al., 2009):
@@ 7.2 Implementation Detail
    interior pixel ($k_{i,j}=\textrm{interior}$), then the weights of all adjacent pixels are set to $1$, i.e., $w_{i+1,j}=w_{i-1,j}=w_{i,j+1}=w_{i,j-1}=1$.
@@
    boundary pixel with Dirichlet condition ($k_{i,j}=\textrm{Dirichlet}$), then Eq. (14) is not executed. Instead, the color of the pixel is determined by the boundary condition.
@@
    boundary pixel with Neumann condition ($k_{i,j}=\textrm{Neumann}$), then the weight of the adjacent pixel $w_{i+1,j}$ is set to $1$ (otherwise 0) if it is non-separated by a patch boundary curve $s^{h}_{i,j}=\textrm{open}$ (analogous for the other neighbor pixels).
@@ 8 Results
In the following section, we present our results using scenes that contain gradient meshes, diffusion curves, and Poisson curves. We made all test scenes available online (Tian and Günther, 2025). All images are rendered at a resolution of $1024\times 1024$.
(a) input meshes only
(b) input curves only
@@ 8.1 Qualitative Results
For all test scenes, we show results of only rasterizing the gradient meshes in Fig. 18(a) and only rasterizing the input curves in (b). This way, we can see which parts of the image are modeled with gradient meshes or diffusion/Poisson curves, respectively. To illustrate the input primitives, we depict the gradient mesh control polygon with solid lines, where different colors are applied for each mesh. Input boundary curves are shown as double-sided curve with an illustration of the respective boundary condition on the left and right. Dirichlet conditions directly show the corresponding color, while Neumann conditions are illustrated with a dash-dot pattern. Poisson curves are visualized with a long dash pattern. In (c), the undirected edge graph is visualized by solid black curves and the vertices are depicted with black circles. Note that Poisson curves are not part of the edge graph, since they are rendered additively onto the target Laplacian function. The unified patch representation in (d) gives all pixels that belong to the same region a common color. The colors are assigned from a custom color table such that the regions are visually distinguishable. The color is not related to the image content. The last column in (e) shows the final result.
@@
Inspired from an ancient Chinese painting (the Auspicious Cranes by Emperor Huizong), we modeled a crane with diffusion curves, which is shown in Fig. 1. The boundary of the crane is largely modeled with Neumann conditions. We modeled the clouds with two $3\times 4$ gradient meshes. In addition, we added diffusion curves and Poisson curves on the gradient meshes for the layered and detailed effects of clouds. The background is formed from a $3\times 3$ gradient mesh that mimics the similar color transition of the inspirational image.
@@
Fig. 18 (first row) shows an extension of the poivron scene from Orzan et al. (2008). We replaced the background with a $30\times 30$ gradient mesh that was fitted with Adobe Illustrator to a photograph inside a restaurant. The gradient mesh’s smoothness creates the impression of defocus blur, putting emphasis on the pepper in the foreground. We took a bite out of the pepper by adding further diffusion curves. The outer boundary of the pepper is modeled with homogeneous Neumann conditions, such that it can be placed in the scene without causing color diffusion into the background. The shadow on the table is modeled with Poisson curves.
@@
Fig. 18 (second row) shows a colorful sunset. The color gradient in the background contains multiple hues and is modeled with a $10\times 3$ gradient mesh. We also deformed the spatial control points for artistic effects. Since it is interrupted by other curve primitives in-between, the color gradient would have been difficult to model with diffusion curves only. The sun is drawn using a diffusion curve, which allows for the modeling of the faint glow into the background. On the top and bottom of the image, we placed diffusion curves that contrast the nature background with expressive wavy shapes.
@@
In Fig. 18 (third row), we modeled two soap bubbles. We used a $1\times 3$ gradient mesh to model smooth color transitions in the sky. The bubbles are created by two $4\times 4$ gradient meshes. The smooth color transitions are difficult to model directly with diffusion curves. This example shows that a vivid scene can be represented by a small number of primitives when combining mesh-based and curve-based methods.
Ladybug
The test scene in Fig. 18 (fourth row) is based on the ladybug diffusion curve model of Orzan et al. (2008). We made minor adjustments to the original curve endpoints to remove edge crossings that would otherwise cause color leakage. In addition, we set the outer boundary of the ladybug to Neumann condition. We placed the ladybug in a beach scene, which contains an ocean and a blue sky. The background is modeled using a $11\times 3$ gradient mesh. In addition, we modeled a leaf on which the ladybug sits with a $6\times 2$ gradient mesh, which captures the subtle color changes without smoothing. The leaf veins are created with Poisson curves, and the outer boundary is modeled with diffusion curves to add scattering.
Portal
The last test scene in Fig. 18 (fifth row) shows a portal. We created the portal with three gradient meshes: a $4\times 3$ background mesh, and two $3\times 4$ gradient meshes for the outer and internal parts of the portal. The rendered portal contains smooth color transitions to represent light and shadows. Inside the portal, We modified the zephyr diffusion curve model of Orzan et al. (2008). The diffusion curves diffuse color into the portal. The mesh Laplacian provides richer effects and more potential for artistic control compared to curve-only results.
(a) reference image (ours)
(b) gradient meshes (manual)
(c) diffusion curves (Orzan)
Orzan et al. (2008)
(d) diffusion curves (isocontours)
Zhao et al. (2017)
Refer to captionRefer to captionRMSE: 0.013
Refer to captionRefer to captionRMSE: 0.018
Refer to captionRefer to captionRMSE: 0.008
Refer to captionRefer to captionRMSE: 0.026
Refer to captionRefer to captionRMSE: 0.089
Refer to captionRefer to captionRMSE: 0.008
Refer to captionRefer to captionRMSE: 0.032
Refer to captionRefer to captionRMSE: 0.033
Refer to captionRefer to captionRMSE: 0.013
Refer to captionRefer to captionRMSE: 0.022
Refer to captionRefer to captionRMSE: 0.024
Refer to captionRefer to captionRMSE: 0.011
Refer to captionRefer to captionRMSE: 0.021
Refer to captionRefer to captionRMSE: 0.043
Refer to captionRefer to captionRMSE: 0.007
Refer to captionRefer to captionRMSE: 0.064
Refer to captionRefer to captionRMSE: 0.022
Refer to captionRefer to captionRMSE: 0.011
Figure 19: Taking our result (a) as a reference, we test how well gradient meshes only (b) and diffusion curves only (c)-(d) approximate our result. Difference images show that gradient meshes often better capture inhomogeneous gradients in the background, while diffusion curves capture details in the foreground.
@@
8.2 Comparison with Vectorizations
So far, there exists no other method to compare with that supports the image synthesis for scenes that contain both gradient meshes and diffusion curves. However, once an automatic vectorization algorithm is developed for our unified formulation, it should be possible to compare how well gradient meshes only, diffusion curves only, and our unified formulation are able to represent the same raster image. For now, we can do a slightly different experiment. In Fig. 19, we take the output from our unified renderer (a) and vectorize this image using gradient meshes (b) and diffusion curves (c)-(d). For each image, the input primitives and a difference image to (a) are shown. The gradient meshes are placed manually using Adobe Illustrator. We can see that gradient meshes have a difficult time adjusting to local structures since they always subdivide entire rows or columns, see for example the swirls behind the crane. Diffusion curves are useful for modeling sharp edges and details, which can be seen in the foreground objects across all test scenes. In (c), the diffusion curves are vectorized using Canny edge detection (Orzan et al., 2007) on the multi-channel color gradient of Di Zenzo (1986), as done by Orzan et al. (2008). In (d), edges are found using isocontours in the residual field as done by Zhao et al. (2017), although without their subsequent shape optimization. For both (c) and (d), we use the vectorization methods to fill the areas of the gradient meshes, i.e., we take our diffusion curves as initial set to start from. We can see in the difference images that many diffusion curves are needed to represent the bi-cubic color gradients of a gradient mesh, which is apparent in scenes with complex backgrounds, such as in the pepper scene. While (d) obtains a low reconstruction error, the high curve density is not necessarily user-friendly. A gradient mesh could have represented the inhomogeneous color gradients more compactly. With our unified approach, both primitives can be combined in the same scene.
@@ 8.3 User Interaction
When individual input primitives are moved in the scene by the user, intersections with other primitives will occur that lead to the creation or removal of patches. To ensure a consistent editing experience, we made sure that all user input is specified on the extended gradient meshes, extended diffusion curves, and Poisson curves alone. On the front end, the user is not accessing the edge graph or the patch representation, which are both computed fully automatically from the input primitives. Fig. 20 shows the rendered output when taking a closed diffusion curve and moving it across a scene. Intersections with other diffusion curves and gradient meshes are created, which causes new patches to form or disappear. Using our data structures, the results remain consistent throughout the editing process.
Table 1: Performance measurements for the various test scenes used in the paper. The columns list the corresponding figure, the number of input diffusion curves (#DC), the number of input Poisson curves (#PC), the number of input gradient meshes (#GM) for which we list the grid resolution in brackets, the number of vertices (#V) and edges (#E) in the edge graph, the number of patches (#P), the computation time to form the edge graph, the computation time to traverse the edge graph, the time needed to resolve the nesting of patches, and the number of Jacobi iterations for rendering. All timings are in milliseconds.
Scene Figure #DCs #PCs #GMs #Vs #Es #Ps Form graph Trav. graph Form patches It.
Crane Fig. 1 68 5 3 160 124 8 16 5 13 1 k
Pepper Fig. 18 (first row) 36 3 1 184 165 8 11 9 12 10 k
Sunset Fig. 18 (second row) 38 6 1 88 108 37 16 2 7 15 k
Bubble Fig. 18 (third row) 9 0 3 62 53 3 2 1 1 15 k
Ladybug Fig. 18 (fourth row) 59 8 2 111 113 20 14 2 16 2 k
Portal Fig. 18 (fifth row) 87 1 3 211 167 17 20 3 14 15 k
@@ 8.4 Performance Analysis
We evaluated our method on an Intel Core i9-10980XE with 4.6 Ghz and an Nvidia RTX 2080 TI GPU. The performance measurements are reported in Table 1. To give context we report for each scene statistics about the complexity of the input data (number of diffusion curves, number of Poisson curves, number of gradient meshes), statistics on the complexity of the edge graph (number of vertices and edges), and the number of patches in our patch representation. We report the time it takes to compute the edge graph, to traverse the edge graph during the patch construction, and to resolve the nesting relationships.
Across all test scenes, the edge graph construction was at around 12-17 ms, the edge graph traversal at 4-10 ms, and the patch creation at 13-45 ms. In sum, we achieved across all scenes an interactive frame rate (29-67 milliseconds per frame). While we incorporated bounding volume hierarchies to accelerate the intersection and containment tests, we think that further improvements could include partial updates of both the edge graph and the patch representation whenever an object moves. Further, a hierarchical discretization could be used to accelerate the winding angle tests. Despite us rebuilding the graph and patches from scratch, we obtained interactive results.
01,0001{,}0002,0002{,}00002,0002{,}0004,0004{,}0006,0006{,}0008,0008{,}00010,00010{,} [^0] Total number of vertices in edge graphTime (ms)
025,00025{,}00050,00050{,}00075,00075{,}000100,000100{,}00002002004004006006008008001,0001{,} [^0] Total number of vertices in polylinesTime (ms)
Figure 21: Scalability experiments (total time in milliseconds) for varying input complexity (left) and discretization (right) on an Intel Core i7-6700K.
To assess the scalability, we randomly generated a varying number of diffusion curves in the unit square. The x-coordinates of the control points were set to $(0,0.3,0.7,1.0)$. The y-coordinates are random numbers in $[0,1]$. This test leads to a large number of edge crossings. In Fig. 21 (left), we plot the total construction time as a function of the number of vertices $|\mathcal{V}|$ in the edge graph, which is a measure of the scene complexity. We observe that the runtime scales quadratically in the number of graph vertices, which can be expected since each edge is tested for intersections against each other edge.
Further, we examine the scalability of the algorithm for varying polyline discretizations in the crane scene. In Fig. 21 (right), we varied the threshold $\tau\in\{0.1,0.05,\dots,0.0001\}$, which leads to an increasing number of polyline vertices. We observe that the total construction time scales nearly linear in the number of polyline vertices. This observation holds for the other scenes, as well, but with a different slope.
@@ 8.5 Discussion
The edge graph and our resulting patch construction directly depend on the quality of the input primitives. If the input primitives contain tiny gaps or small crossings, then a corresponding patch representation would carry on those artifacts. In recent years, more exact vector graphics rendering algorithms have been developed and utilized, such as Walk-on-Spheres (Sawhney and Crane, 2020), Walk-on-Boundaries (Sugimoto et al., 2023), or Walk-on-Stars (Sawhney et al., 2023), or Walkin’ Robin (Miller et al., 2024b). These approaches likewise suffer from inaccuracies in the scene description, since they are able to resolve the color transport through narrow gaps. In Section 5.3, we described our scene preprocessing, which closes gaps to prevent color leaking. To alleviate the need for such preprocessing, more research towards user interfaces that naturally steer the user towards more accurate scene descriptions would be helpful for both scene modeling and rendering.
@@ 9 Conclusions
In this paper, we unified the mathematical modeling of mesh-based vector graphics and curve-based vector graphics, which allows for the first time to include both types of primitives in the same scene. For this, we rephrased the interpolation inside gradient meshes as solution to a Poisson problem by searching for a color field that matches the Laplacian of the gradient mesh. We developed a four-stage pipeline that takes gradient meshes, diffusion curves, and Poisson curves as input. First, we extended gradient meshes and diffusion curves by enabling the specification of Neumann conditions in addition to the usual Dirichlet conditions. Second, our approach handles arbitrary intersections of diffusion curves and gradient meshes, since all intersections are resolved through the construction of an edge graph. Third, from the edge graph, non-overlapping patches with well-defined boundary conditions and a target Laplacian are derived. Fourth, a standard Poisson problem is solved in the interior of each patch to synthesize the output image. The final rasterized images can be computed with any off-the-shelf Poisson solver.
In the past, research on mesh-based and curve-based smooth vector graphics followed independent research threads, which concentrated on the editing, rasterization, or vectorization of the individual primitives. We hope that the unified treatment will spur further research that promotes synergies of the two approaches, for example regarding novel image synthesis or vectorization methods. The next step in the generalization will include the incorporation of bi-harmonic diffusion curves Finch et al. (2011), for which harmonic solutions can be searched in the Laplacian domain. Similar to common content creation tools, it could be useful to introduce multiple layers of smooth vector graphics. Further, it would be interesting to develop a vectorization algorithm for the unified patch description. It would be imaginable to divide the input domain into regions. Per region, the algorithm could decide whether it is better to fit a gradient mesh or a set of diffusion curves. We leave the development of a vectorization algorithm to future work.